# Assimilative Causal Inference

### v1

@@ A.5 Instantaneous Discovery of the Causal Influence Range (CIR)
The divergence between these distributions reveals the temporal decay of $\mathbf{y}(t)$’s causal influence on future $\mathbf{x}$ values as $T^{\prime}$ increases from $t$ towards $T$.
@@ A.5.1 Quantifying the CIR
$\delta(T^{\prime};t):=\mathcal{P}\big{(}p_{t}(\mathbf{y}|\mathbf{x}(s\leq T)),p_{t}(\mathbf{y}|\mathbf{x}(s\leq T^{\prime}))\big{)},\quad 0\leq t\leq T^{\prime}\leq T$ (1.16)
@@
$\hat{\delta}(\tau;t):=\delta(t+\tau(T-t);t),\quad\tau\in\mathrm{I}.$ (1.17)
@@
For chaotic systems, $\mathcal{M}(t)$ typically occurs at $\tau=0$ or shortly thereafter. The interval where $\hat{\delta}(\tau;t)$ remains large suggests significant causal influence under ACI, since it indicates that there is substantial information gain to be incurred by additionally incorporating the future observations of $\mathbf{x}$ from $[h(\tau),T]$.
@@
$\widetilde{\mathrm{CIR}}_{\mathbf{y}(t)\rightarrow\mathbf{x}}(\varepsilon):=[t,t+\tilde{\tau}_{\mathbf{y}(t)\rightarrow\mathbf{x}}(\varepsilon)].$ (1.20)
As aforementioned, for most practical chaotic dynamical systems, the maximum information deficit $\mathcal{M}(t)$ typically occurs at $\tau=0$ or shortly thereafter. This reflects the characteristic memory decay in such systems, where the influence on the state estimation of $\mathbf{y}(t)$ from the future observations of $\mathbf{x}$ in $[h(\tau),T]$ diminishes with increasing lead time $\tau$. Note also that under the mild assumption that $\hat{\delta}(\cdot;t)$ is Lebesgue measurable for each $t\in[0,T]$, which is a condition satisfied when $\delta(\cdot;t)$ is Borel measurable (that holds for the relative entropy and most complex turbulent systems found in practice), we obtain the fundamental inequality:
@@ A.5.2 Subjective and Objective Perspectives of the CIR
  • Subjective CIR: Denoted by $\tilde{\tau}_{\mathbf{y}(t)\rightarrow\mathbf{x}}(\varepsilon)$ (length) and $\widetilde{\mathrm{CIR}}_{\mathbf{y}(t)\rightarrow\mathbf{x}}(\varepsilon)$ (interval), marked with tildes to emphasize their $\varepsilon$-dependence.
@@
[…]
where $\mathcal{M}(t)=\max_{\tau\in[0,1]}\hat{\delta}(\tau;t)$. Dividing $\mathcal{M}(t)$ guarantees the unit of the objective CIR is “time”. This averaging process: (a) yields dimensionally consistent results (time units), (b) guarantees $\tau_{\mathbf{y}(t)\rightarrow\mathbf{x}}\in[0,T-t]$, and (c) defaults to 00 when $\mathcal{M}(t)=0$ (no causal link).
@@
$\mathrm{CIR}_{\mathbf{y}(t)\rightarrow\mathbf{x}}:=[t,t+\tau_{\mathbf{y}(t)\rightarrow\mathbf{x}}],\quad t\in[0,T].$ (1.23)
By construction, this objective interval is always contained within the maximal subjective CIR.
@@ A.5.3 Efficient Computation of the Objective CIR
While (1.22) offers a theoretically sound definition of the objective CIR length, its practical implementation faces significant computational challenges. The subjective CIR length $\tilde{\tau}_{\mathbf{y}(t)\rightarrow\mathbf{x}}(\varepsilon)$ generally admits no analytical solution across the full parameter space $\varepsilon\in[0,\mathcal{M}(t)]$ and $t\in[0,T]$. Numerically evaluating the integral in (1.22) requires repeated smoother computations and leads to a computational complexity that scales at best quadratically with the number of discretization points adopted for quadrature.
@@
Assume that $\exists\mathcal{M}(t):=\lVert\hat{\delta}(\cdot;t)\rVert_{L^{\infty}(\mathrm{I})}=\lVert\delta(\cdot;t)\rVert_{L^{\infty}(\mathrm{[t,T]})}>0$ for each $t\in[0,T]$. Then, for $t\in[0,T]$:
@@
[…]
@@
First, observe that $\lambda_{\mathrm{I}}(\cdot)$ defines a probability measure on $\mathrm{I}=[0,1]$. Under the Lebesgue measurability assumption for $\hat{\delta}(\cdot;t)$, we can then interpret $\hat{\delta}(\cdot;t)$ as a continuous random variable on the probability space $(\mathrm{I},\mathcal{L}_{\mathrm{I}},\lambda_{\mathrm{I}})$, $\mathcal{L}_{\mathrm{I}}$ as the $\sigma$-algebra of Lebesgue measurable subsets of $\mathrm{I}$, where the support of $\hat{\delta}(\cdot;t)$ is $[0,\mathcal{M}(t)]$.
@@
The theorem’s main result follows by dividing through by $\mathcal{M}(t)>0$, multiplying by $T-t$, and applying the inequality in (1.21). The equality condition follows immediately from (1.21), since the equality there holds if and only if $\hat{\delta}(\cdot;t)$ is nonincreasing on $\mathrm{I}$. ∎
@@
$\uptau^{\text{approx}}_{\mathbf{y}(t)\rightarrow\mathbf{x}}:=\frac{T-t}{\mathcal{M}(t)}\int^{1}_{0}\hat{\delta}(\tau;t)\mathrm{d}\tau=\frac{1}{\mathcal{M}(t)}\int^{T}_{t}\delta(T^{\prime};t)\mathrm{d}T^{\prime}\leq\uptau_{\mathbf{y}(t)\rightarrow\mathbf{x}},\quad t\in[0,T].$ (1.25)
The inequality reduces to an equality precisely when $\hat{\delta}(\cdot;t)$ is nonincreasing on $\mathrm{I}$, which is a condition satisfied in most practical applications. A dimensional analysis of (1.25) reveals that $\tau^{\text{approx}}_{\mathbf{y}(t)\rightarrow\mathbf{x}}$ maintains consistent time units, confirming the physical validity of this approximation. This consistency mirrors the dimensional properties established for the exact objective CIR length in (1.22).
@@ B.1 Conditional Gaussian Nonlinear Systems (CGNS)
Numerical approximations, such as ensemble or particle methods, have to be adopted to find the filter and smoother solutions for general nonlinear dynamics systems. Nevertheless, analytical solutions are available for the data assimilation solutions for a broad class of nonlinear systems, known as conditional Gaussian nonlinear systems (CGNSs) [15].
@@ B.1.1 The Modeling Framework
A CGNS consists of two Itô diffusion processes and has the following form [15]:
@@
where the matrices $\boldsymbol{\Lambda}^{\mathbf{x}}$, $\boldsymbol{\Lambda}^{\mathbf{y}}$, $\boldsymbol{\Sigma}_{1}^{\mathbf{x}}$, $\boldsymbol{\Sigma}_{2}^{\mathbf{x}}$, $\boldsymbol{\Sigma}_{1}^{\mathbf{y}}$, $\boldsymbol{\Sigma}_{2}^{\mathbf{y}}$ and the vectors $\mathbf{f}^{\mathbf{x}}$, $\mathbf{f}^{\mathbf{y}}$ can contain arbitrarily nonlinear functions of $\mathbf{x}$. The state variable $\mathbf{y}$ appears in the system in a conditionally linear way but it can interact nonlinearly with $\mathbf{x}$. Therefore, the system is overall highly nonlinear and can generate strong non-Gaussian feature in both the marginal and joint distributions, corresponding to extreme events, intermittency, regime switching, etc. When all these matrices and vectors become only a function of time, i.e., they do not depend on $\mathbf{x}$, then the filtering and smoothing of (2.1) collapses to the setup of the classical Kalman-Bucy filter [11] and Rauch-Tung-Striebel smoother [20].
The CGNS framework broadly applies to many complex nonlinear systems across disciplines, including physics-constrained stochastic models (e.g., noisy Lorenz systems, low-order Charney-DeVore flows, and topographic mean-flow interaction paradigms), stochastically coupled reaction-diffusion systems like FitzHugh-Nagumo neural models and SIR epidemic models, and multiscale geophysical flow models such as stochastic Boussinesq equations and forced rotating shallow water equations. This framework has also been successfully adapted to model realistic phenomena like the Madden-Julian oscillation and Arctic sea ice dynamics. See a gallery of examples in [5]. On the other hand, many multiscale systems with general nonlinearities can be effectively approximated as CGNSs with minimal error. In geophysical and fluid systems, where nonlinearities often take quadratic forms, this approximation preserves all nonlinear interactions involving the large-scale variables $\mathbf{x}$ and their couplings with the small-scale fluctuations in $\mathbf{y}$. Only the self-interactions of $\mathbf{y}$ are replaced by effective noise and damping terms, which is an approach rigorously justified when $\mathbf{y}$ represents fast, unresolved scales. This strategy aligns with the stochastic mode reduction framework developed by Majda, Timofeyev, and Vanden-Eijnden (the MTV method) [17].
@@ B.1.2 Filtering and Smoothing Distributions for CGNSs
Theorem B.1 (Optimal nonlinear filter state estimation equations for CGNSs [15]).
@@
$\boldsymbol{\mu}_{\text{{f}}}(t):=\mathbb{E}\left[\mathbf{y}(t)\big{|}\mathbf{x}(s\leq t)\right]\quad\text{and}\quad\mathbf{R}_{\text{{f}}}(t):=\mathbb{E}\left[(\mathbf{y}(t)-\boldsymbol{\mu}_{\text{{f}}}(t))(\mathbf{y}(t)-\boldsymbol{\mu}_{\text{{f}}}(t))^{\mathtt{T}}\big{|}\mathbf{x}(s\leq t)\right],$
@@
Theorem B.2 (Optimal nonlinear smoother state estimation backward equations for CGNSs [15]).
@@
$\boldsymbol{\mu}_{\text{{s}}}(t):=\mathbb{E}\left[\mathbf{y}(t)\big{|}\mathbf{x}(s\leq T)\right]\quad\text{and}\quad\mathbf{R}_{\text{{s}}}(t):=\mathbb{E}\left[(\mathbf{y}(t)-\boldsymbol{\mu}_{\text{{s}}}(t))(\mathbf{y}(t)-\boldsymbol{\mu}_{\text{{s}}}(t))^{\mathtt{T}}\big{|}\mathbf{x}(s\leq T)\right],$
@@ B.2 Online Smoother for CGNSs
The theoretical framework in Section AA.5 requires computation of $\delta(T^{\prime};t)$ in (1.16) for all $t\in[0,T]$ and $T^{\prime}\in[t,T]$ to evaluate either the subjective or objective CIR of an (conditional) assimilative causal link. For CGNSs, this calculation is now feasible thanks to recent advances in online smoothing algorithms [2], which allow real-time computation of smoother distributions as $\mathbf{x}$ observations arrive sequentially.
Let $\mathcal{T}(\Delta t)=\{t_{j}\}_{j\in\llbracket N\rrbracket}$ be a uniform partition of $[0,T]$ with mesh size $0<\Delta t\ll 1$ and $t_{j}=j\Delta t$, for $j\in\llbracket N\rrbracket:=\{0,1,\ldots,N\}$, where $N=T/\Delta t$. In what follows we use a superscript notation $\cdot^{\,j}$ to denote the discrete approximation to the continuous form of the respective function when evaluated on $t_{j}$, e.g., $\boldsymbol{\Lambda}^{\mathbf{x},j}:=\boldsymbol{\Lambda}^{\mathbf{x}}(t_{j},\mathbf{x}(t_{j}))$ and $\mathbf{x}^{j}:=\mathbf{x}(t_{j})$. Similar to the continuous-time setup, $\big{(}\cdot\big{|}\mathbf{x}(s\leq n)\big{)}$ denotes that we are conditioning on the $\sigma$-algebra generated by $\{\mathbf{x}^{0},\ldots,\mathbf{x}^{n}\}$. Define the following auxiliary matrices for $j\in\llbracket N\rrbracket$:
@@
Theorem B.3 (Optimal online smoother for CGNSs [2]).
@@
for $n\in\llbracket N\rrbracket$ and $j=n$ are given by $\boldsymbol{\mu}_{\text{{s}}}^{n,n}=\boldsymbol{\mu}_{\text{{f}}}^{n}$ and $\mathbf{R}_{\text{{s}}}^{n,n}=\mathbf{R}_{\text{{f}}}^{n}$, since the smoother and filter Gaussian statistics coincide at the current end point $t_{n}$. Then, for $n\in\{1,\ldots,N\}$ and $j=n-1$ they are instead given by the following equations:
@@
$\mathbf{R}_{\text{{s}}}^{j,n}$ $=\mathbf{R}_{\text{{s}}}^{j,n-1}+\mathbf{D}^{j,n-2}\left(\mathbf{R}_{\text{{s}}}^{n-1,n}-\mathbf{R}_{\text{{f}}}^{n-1}\right)(\mathbf{D}^{j,n-1})^{\mathtt{T}},$ (2.6b)
@@
$\mathbf{D}^{n-1,n-2}:=\mathbf{I}_{l\times l}\quad\&\quad\mathbf{D}^{j,n-2}:=\overset{{\curvearrowright}}{\prod^{n-2}_{i=j}}\mathbf{E}^{i}:=\mathbf{E}^{j}\mathbf{E}^{j+1}\cdots\mathbf{E}^{n-2}.$ (2.7)

### v2

@@ A.5 Dynamic Discovery of the Causal Influence Range (CIR)
The divergence between these distributions reveals the temporal decay of $\mathbf{y}(t)$’s causal influence on future $\mathbf{x}$ values as $T^{\prime}$ increases from $t$ toward $T$.
@@ A.5.1 Quantifying the CIR
$\delta(T^{\prime};t):=\mathcal{P}\big(p_{t}(\mathbf{y}|\mathbf{x}(s\leq T)),p_{t}(\mathbf{y}|\mathbf{x}(s\leq T^{\prime}))\big),\qquad 0\leq t\leq T^{\prime}\leq T$ (1.16)
@@
$\hat{\delta}(\tau;t):=\delta(t+\tau(T-t);t),\qquad\tau\in\mathrm{I}.$ (1.17)
@@
For chaotic systems, $\mathcal{M}(t)$ typically occurs at $\tau=0$ or shortly thereafter. The interval where $\hat{\delta}(\tau;t)$ remains large suggests significant causal influence under ACI, since it indicates that there is substantial information gain to be incurred by additionally incorporating the future observations of $\mathbf{x}$ from $h(\tau),T]$.
@@
$\widetilde{\mathrm{CIR}}_{\mathbf{y}(t)\rightarrow\mathbf{x}}(\varepsilon):=[t,t+\overset{\sim}{\uptau}_{\mathbf{y}(t)\rightarrow\mathbf{x}}(\varepsilon)].$ (1.20)
As aforementioned, for most practical chaotic dynamical systems, the maximum information deficit $\mathcal{M}(t)$ typically occurs at $\tau=0$ or shortly thereafter. This reflects the characteristic memory decay in such systems, where the influence on the state estimation of $\mathbf{y}(t)$ from the future observations of $\mathbf{x}$ in $h(\tau),T]$ diminishes with increasing lead time $\tau$. Note also that under the mild assumption that $\hat{\delta}(\cdot;t)$ is Lebesgue measurable for each $t\in[0,T]$, which is a condition satisfied when $\delta(\cdot;t)$ is Borel measurable (that holds for the relative entropy and most complex turbulent systems found in practice), we obtain the fundamental inequality:
@@ A.5.2 Subjective and Objective Perspectives of the CIR
  • Subjective CIR: Denoted by $\overset{\sim}{\uptau}_{\mathbf{y}(t)\rightarrow\mathbf{x}}(\varepsilon)$ (length) and $\widetilde{\mathrm{CIR}}_{\mathbf{y}(t)\rightarrow\mathbf{x}}(\varepsilon)$ (interval), marked with tildes to emphasize their $\varepsilon$-dependence.
@@
$\uptau_{\mathbf{y}(t)\rightarrow\mathbf{x}}:=\frac{1}{\mathcal{M}(t)}\int_{0}^{\mathcal{M}(t)}\overset{\sim}{\uptau}_{\mathbf{y}(t)\rightarrow\mathbf{x}}(\varepsilon){\mathrm{d}}\varepsilon,\qquad t\in[0,T],$ (1.22)
where $\mathcal{M}(t)=\max_{\tau\in[0,1]}\hat{\delta}(\tau;t)$. Dividing by $\mathcal{M}(t)$ guarantees the unit of the objective CIR is temporal. This averaging process: (a) yields dimensionally consistent results (time units), (b) guarantees $\uptau_{\mathbf{y}(t)\rightarrow\mathbf{x}}\in[0,T-t]$, and (c) defaults to 0 when $\mathcal{M}(t)=0$ (no causal link).
@@
$\mathrm{CIR}_{\mathbf{y}(t)\rightarrow\mathbf{x}}:=[t,t+\uptau_{\mathbf{y}(t)\rightarrow\mathbf{x}}],\qquad t\in[0,T].$ (1.23)
By construction, this objective interval is always contained within the maximal subjective CIR and always contains the minimal one.
@@ A.5.3 Efficient Computation of the Objective CIR
While (1.22) offers a theoretically sound definition of the objective CIR length, its practical implementation faces significant computational challenges. The subjective CIR length $\overset{\sim}{\uptau}_{\mathbf{y}(t)\rightarrow\mathbf{x}}(\varepsilon)$ generally admits no analytical solution across the full parameter space $\varepsilon\in[0,\mathcal{M}(t)]$ and $t\in[0,T]$. Numerically evaluating the integral in (1.22) requires repeated smoother computations and leads to a computational complexity that scales at best quadratically with the number of discretization points adopted for quadrature.
@@
Assume that there exists $\mathcal{M}(t):=\lVert\hat{\delta}(\cdot;t)\rVert_{L^{\infty}(\mathrm{I})}=\lVert\delta(\cdot;t)\rVert_{L^{\infty}(\mathrm{[t,T]})}>0$ for each $t\in[0,T]$. Then, for $t\in[0,T]$:
@@
$\leq\frac{1}{\mathcal{M}(t)}\int_{0}^{\mathcal{M}(t)}\overset{\sim}{\uptau}_{\mathbf{y}(t)\overset{{\text{{ACI}}}}{\longrightarrow}\mathbf{x}}(\varepsilon)\mathrm{d}\varepsilon=\uptau_{\mathbf{y}(t)\overset{{\text{{ACI}}}}{\longrightarrow}\mathbf{x}},$
where $\hat{\delta}(\tau;t)$, $\mathrm{J}_{t}(\varepsilon)$, and $\overset{\sim}{\uptau}_{\mathbf{y}(t)\overset{{\text{{ACI}}}}{\longrightarrow}\mathbf{x}}(\varepsilon)$ are defined in (1.17), (1.18), and (1.19), respectively. The equality (in the last inequality in (1.24)) is achieved if and only if $\hat{\delta}(\cdot;t)$ is a nonincreasing function in $\mathrm{I}$. In that case, we also have:
@@
First, observe that $\lambda_{\mathrm{I}}(\cdot)$ defines a probability measure on $\mathrm{I}=[0,1]$. Under the Lebesgue measurability assumption for $\hat{\delta}(\cdot;t)$, we can then interpret $\hat{\delta}(\cdot;t)$ as a continuous random variable on the probability space $(\mathrm{I},\mathcal{L}_{\mathrm{I}},\lambda_{\mathrm{I}})$, $\mathcal{L}_{\mathrm{I}}$ as the $\sigma$-algebra of Lebesgue measurable subsets of $\mathrm{I}$, where the support of $\hat{\delta}(\cdot;t)$ is $0,\mathcal{M}(t)]$.
@@
The theorem’s main result follows by dividing through by $\mathcal{M}(t)>0$, multiplying by TtT-t, and applying the inequality in (1.21). The equality condition follows immediately from (1.21), since the equality there holds if and only if $\hat{\delta}(\cdot;t)$ is nonincreasing on $\mathrm{I}$. ∎
@@
$\uptau^{\text{approx}}_{\mathbf{y}(t)\rightarrow\mathbf{x}}:=\frac{T-t}{\mathcal{M}(t)}\int^{1}_{0}\hat{\delta}(\tau;t)\mathrm{d}\tau=\frac{1}{\mathcal{M}(t)}\int^{T}_{t}\delta(T^{\prime};t)\mathrm{d}T^{\prime}\leq\uptau_{\mathbf{y}(t)\rightarrow\mathbf{x}},\qquad t\in[0,T].$ (1.25)
The inequality reduces to an equality precisely when $\hat{\delta}(\cdot;t)$ is nonincreasing on $\mathrm{I}$, which is a condition satisfied in most practical applications. A dimensional analysis of (1.25) reveals that $\uptau^{\text{approx}}_{\mathbf{y}(t)\rightarrow\mathbf{x}}$ maintains consistent time units, confirming the physical validity of this approximation. This consistency mirrors the dimensional properties established for the exact objective CIR length in (1.22).
@@ B.1 Conditional Gaussian Nonlinear Systems (CGNS)
Numerical approximations, such as ensemble or particle methods, have to be adopted to find the filter and smoother solutions for general nonlinear dynamics systems. Nevertheless, analytical solutions are available for the data assimilation solutions for a broad class of nonlinear systems, known as conditional Gaussian nonlinear systems (CGNSs) [liptser2001statistics].
@@ B.1.1 The Modeling Framework
A CGNS consists of two Itô diffusion processes and has the following form [liptser2001statistics]:
@@
where the matrices $\boldsymbol{\Lambda}^{\mathbf{x}}$, $\boldsymbol{\Lambda}^{\mathbf{y}}$, $\boldsymbol{\Sigma}_{1}^{\mathbf{x}}$, $\boldsymbol{\Sigma}_{2}^{\mathbf{x}}$, $\boldsymbol{\Sigma}_{1}^{\mathbf{y}}$, $\boldsymbol{\Sigma}_{2}^{\mathbf{y}}$ and the vectors $\mathbf{f}^{\mathbf{x}}$, $\mathbf{f}^{\mathbf{y}}$ can contain arbitrarily nonlinear functions of $\mathbf{x}$. The state variable $\mathbf{y}$ appears in the system in a conditionally linear way but it can interact nonlinearly with $\mathbf{x}$. Therefore, the system is overall highly nonlinear and can generate strong non-Gaussian feature in both the marginal and joint distributions, corresponding to extreme events, intermittency, regime switching, etc. When all these matrices and vectors become only a function of time, i.e., they do not depend on $\mathbf{x}$, then the filtering and smoothing of (2.1) collapses to the setup of the classical Kalman-Bucy filter [kalman1961new] and Rauch-Tung-Striebel smoother [rauch1965maximum].
The CGNS framework broadly applies to many complex nonlinear systems across disciplines, including physics-constrained stochastic models (e.g., noisy Lorenz systems, low-order Charney-DeVore flows, and topographic mean-flow interaction paradigms), stochastically coupled reaction-diffusion systems like FitzHugh-Nagumo neural models and SIR epidemic models, and multiscale geophysical flow models such as stochastic Boussinesq equations and forced rotating shallow water equations. This framework has also been successfully adapted to model realistic phenomena like the Madden-Julian oscillation and Arctic sea ice dynamics. See a gallery of examples in [chen2018conditional]. On the other hand, many multiscale systems with general nonlinearities can be effectively approximated as CGNSs with minimal error. In geophysical and fluid systems, where nonlinearities often take quadratic forms, this approximation preserves all nonlinear interactions involving the large-scale variables $\mathbf{x}$ and their couplings with the small-scale fluctuations in $\mathbf{y}$. Only the self-interactions of $\mathbf{y}$ are replaced by effective noise and damping terms, which is an approach rigorously justified when $\mathbf{y}$ represents fast, unresolved scales. This strategy aligns with the stochastic mode reduction framework developed by Majda, Timofeyev, and Vanden-Eijnden (the MTV method) [majda2003systematic].
@@ B.1.2 Filtering and Smoothing Distributions for CGNSs
Theorem B.1 (Optimal nonlinear filter state estimation equations for CGNSs [liptser2001statistics]).
@@
$\boldsymbol{\mu}_{\text{{f}}}(t):=\mathbb{E}\left[\mathbf{y}(t)\big|\mathbf{x}(s\leq t)\right]\qquad\text{and}\qquad\mathbf{R}_{\text{{f}}}(t):=\mathbb{E}\left[(\mathbf{y}(t)-\boldsymbol{\mu}_{\text{{f}}}(t))(\mathbf{y}(t)-\boldsymbol{\mu}_{\text{{f}}}(t))^{\mathtt{T}}\big|\mathbf{x}(s\leq t)\right],$
@@
Theorem B.2 (Optimal nonlinear smoother state estimation backward equations for CGNSs [liptser2001statistics]).
@@
$\boldsymbol{\mu}_{\text{{s}}}(t):=\mathbb{E}\left[\mathbf{y}(t)\big|\mathbf{x}(s\leq T)\right]\qquad\text{and}\qquad\mathbf{R}_{\text{{s}}}(t):=\mathbb{E}\left[(\mathbf{y}(t)-\boldsymbol{\mu}_{\text{{s}}}(t))(\mathbf{y}(t)-\boldsymbol{\mu}_{\text{{s}}}(t))^{\mathtt{T}}\big|\mathbf{x}(s\leq T)\right],$
@@ B.2 Online Smoother for CGNSs
The theoretical framework in Section A.5 requires computation of $\delta(T^{\prime};t)$ in (1.16) for all $t\in[0,T]$ and $T^{\prime}\in[t,T]$ to evaluate either the subjective or objective CIR of an (conditional) assimilative causal link. For CGNSs, this calculation is now feasible thanks to recent advances in online smoothing algorithms [andreou2024adaptive], which allow real-time computation of smoother distributions as $\mathbf{x}$ observations arrive sequentially.
Let $\{t_{j}\}_{j\in\llbracket N\rrbracket}$ be a uniform partition of $[0,T]$ with mesh size $0<\Delta t\ll 1$ and $t_{j}=j\Delta t$, for $j\in\llbracket N\rrbracket:=\{0,1,\ldots,N\}$, where $N=T/\Delta t$. The assumption of uniform $\Delta t$ is without loss of generality [andreou2024adaptive]. In what follows we use a superscript notation $\cdot^{\,j}$ to denote the discrete approximation to the continuous form of the respective function when evaluated on $t_{j}$, e.g., $\boldsymbol{\Lambda}^{\mathbf{x},j}:=\boldsymbol{\Lambda}^{\mathbf{x}}(t_{j},\mathbf{x}(t_{j}))$ and $\mathbf{x}^{j}:=\mathbf{x}(t_{j})$. Similar to the continuous-time setup, $\big(\cdot\big|\mathbf{x}(s\leq n)\big)$ denotes that we are conditioning on the $\sigma$-algebra generated by $\{\mathbf{x}^{0},\ldots,\mathbf{x}^{n}\}$. Define the following auxiliary matrices for $j\in\llbracket N\rrbracket$:
@@
Theorem B.3 (Optimal online smoother for CGNSs [andreou2024adaptive]).
@@
for $n\in\llbracket N\rrbracket$ and $j=n$ are given by $\boldsymbol{\mu}_{\text{{s}}}^{n,n}=\boldsymbol{\mu}_{\text{{f}}}^{n}$ and $\mathbf{R}_{\text{{s}}}^{n,n}=\mathbf{R}_{\text{{f}}}^{n}$, since the smoother and filter Gaussian statistics coincide at the current end point $t_{n}$. Then, for $n\in\{1,\ldots,N\}$ and j=n1j=n-1 they are instead given by the following equations:
@@
$\mathbf{R}_{\text{{s}}}^{j,n}$ $=\mathbf{R}_{\text{{s}}}^{j,n-1}+\mathbf{D}^{j,n-2}\left(\mathbf{R}_{\text{{s}}}^{n-1,n}-\mathbf{R}_{\text{{f}}}^{n-1}\right)(\mathbf{D}^{j,n-2})^{\mathtt{T}},$ (2.6b)
@@
$\mathbf{D}^{n-1,n-2}:=\mathbf{I}_{l\times l}\qquad\&\qquad\mathbf{D}^{j,n-2}:=\overset{{\curvearrowright}}{\prod^{n-2}_{i=j}}\mathbf{E}^{i}:=\mathbf{E}^{j}\mathbf{E}^{j+1}\cdots\mathbf{E}^{n-2}.$ (2.7)